Ticket Number: HEA-NOC/20030407-03 Ticket Status: CLOSE
Ticket Type: Scheduled Ticket Source: HEA-NOC
Ticket Scope: Site Site/Line: UCC
Ticket Owner: [log in to unmask] Problem Fixer: UCC Computer Services
Ticket Opened: Problem Start: 20030409 08:30 UTC +1
Ticket Closed: 20030409 14:15:39 UTC Problem End: 20030409 09:00 UTC +1
UCC upgrading their core switches.
UCC will lose external connectivity and there will be some disruption to their internal network. HEAnet's Cork PoP will not be affected.
20030409 15:15 BN
Upgrade went to plan, ticket being closed.
Time to Fix: